# When Large Language Models Meet UAVs: How Far Are We?

### v1

@@
Abstract.
The integration of unmanned aerial vehicles (UAVs) and large language models (LLMs) has emerged as a research direction of growing interest, with the potential to address challenges in autonomous decision-making, human-UAV interaction, and real-time adaptability. However, existing studies have remained largely in preliminary exploration with a limited understanding of real-world practice, risking a misalignment between academic research and practical needs and hindering the translation of results. To examine and address these potential challenges, we conducted an empirical study of 74 selected papers and 56 public GitHub projects, identified nine task types for LLMs in UAV systems, and quantified their distribution. Our findings show that academic research emphasizes theoretical modeling and task optimization with dispersed attention across tasks. In contrast, industrial projects focus on flight control, task planning, and human-machine interaction, prioritizing operability and efficiency. To further capture industry perspectives, we distributed an online questionnaire. We obtained 52 valid responses: 40.4% of practitioners have attempted to apply LLMs to UAV tasks. We further identify factors that impede real-world integration, including technological maturity, performance, safety, cost, and other considerations. Finally, we highlight challenges for future development and provide recommendations.
@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: CSURjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Software and its engineering Software system structuresccs: Software and its engineering Designing softwareccs: Computer systems organization Embedded and cyber-physical systems
[…]
1. Introduction
In recent years, unmanned aerial vehicles (UAVs) have been rapidly applied in agriculture, disaster response, security inspection, logistics, and transportation, offering significant efficiency gains due to their mobility, low deployment cost, and diverse capabilities (Shakhatreh et al., 2019; Liu et al., 2021; Colomina and Molina, 2014). However, while UAV operations involve complex processes of environmental perception, flight control, and task planning, current UAV systems still primarily rely on manual operation or rule-based pipelines (Gupta et al., 2016), which limits their flexibility and adaptability in dynamic environments and diverse tasks. Consequently, it is still difficult to achieve scalable high-autonomy operations and thereby constraining the broader deployment of UAVs in more complex scenarios (Colomina and Molina, 2014).
Advances in large language models (LLMs) in language understanding, reasoning, and generation (Chang et al., 2024; Zhao et al., 2023a) offer new possibilities to address the above limitations. By enabling UAVs to incorporate reasoning and interpretation in complex environments, LLMs reduce reliance on manual operation and rule-based pipelines, allowing tasks to be decomposed and decisions to be made with greater flexibility (Tian et al., 2025; Wang et al., 2024a; Zeng et al., 2023a; Ahn et al., 2022; Brohan et al., 2023). This capability can significantly enhance the autonomy and adaptability of UAVs, making the integration of UAVs with LLMs an increasingly prominent research direction.
Although existing studies (Javaid et al., 2024; Tian et al., 2025; TAZIR et al., 2023) demonstrate the feasibility of this direction, most remain confined to simulation or controlled environments, with limited coverage of scenarios and task diversity. They often give insufficient consideration to critical factors in real-world applications, such as computational resources, latency, reliability, and compliance (Zhang et al., 2023; Wu et al., 2024), and the defined task metrics do not readily correspond to the complexity of practical demands. These limitations risk creating a mismatch between academic research and actual needs, thereby reducing the practical value of research studies in translating them into technologies usable by industry.
@@ 1. Introduction
RQ1. What is the current research status and focus of large language models in the field of UAVs?
@@
Identifying the specific tasks that LLMs can perform in UAV systems is the first step toward understanding the advantages and limitations of their integration. To this end, we conducted a systematic literature review (SLR) (Kitchenham et al., 2009) and analyzed a total of 997 research papers related to LLMs and UAVs, applying a card sorting method (Spencer, 2009) to summarize and classify their contents. Based on the SLR process, we classified the UAV-LLM integration into four core stages in the UAV workflow—Information Input, Task Planning, Control Execution, and Result Feedback—each corresponding to a set of LLM-enabled tasks. Among these research directions, the top three most frequently studied directions are Single-UAV Task Reasoning and Planning, Control Logic and Command Generation, and Natural Language Command Parsing. Our characterization of UAV-LLM research clarifies how LLMs are currently applied within UAV systems, which also enables us to compare academic research with industrial practices in subsequent analysis.
@@
RQ2. What are the differences between academia and industry in integrating UAVs with large language models?
@@
Due to differences in objectives, resources, and evaluation standards, academia and industry often emphasize different aspects when using new technologies. In this study, we aim to compare these differences and explore how the two sides can be better connected, so that research results are more practical and industrial needs can be more clearly reflected in future research. We collected 1,509 UAV-LLM-related projects from GitHub and performed an empirical analysis of their characteristics and development practices. Our analysis revealed different efforts between academia and industry in the UAV-LLM field. Academic research focuses more on the design of complex system integration and theoretical mechanisms. It is committed to exploring how to use LLMs to improve system autonomy, flexibility, and multi-tasking capabilities. In contrast, industrial practice emphasizes system stability, low-cost implementation, and rapid deployment, and generally tends to adopt efficient solutions. This, to some extent, limits the in-depth application of LLMs in scenarios requiring high timeliness or operating under resource constraints. By revealing the gap between academia and industry, our results could provide practical implications for facilitating real-world UAV-LLM practices.
@@
RQ3. How do real-world practitioners perceive these differences, and what factors contribute to them?
@@
Although our comparative analysis reveals the task preferences of UAV-LLM in different scenarios, the root causes of these differences remain unclear. To this end, we designed a survey for real-world developers, focusing on the key obstacles and design choices faced by practitioners in UAV-LLM integration. The analysis of 52 real-world responses shows that the integration of LLMs with UAVs remains at an early stage. Only 19.05% of teams have deployed such projects, while 59.6% have not attempted integration, citing reasons such as existing technology has met the project requirements (54.8%), insufficient performance of current LLMs (51.6%), and security risks of large language models (45.2%). The assessed technical maturity of all tasks is below three on a 5-point scale, with relatively higher scores for Information Summarization and Report Generation and LLM-based Chatbot, but lower scores for Multi-UAV Task Reasoning and Planning and Control Logic and Command Generation. Furthermore, 82.7% of practitioners prefer Onboard Control (Internal) Modes over direct control, reflecting caution toward applying LLMs in safety-critical operations. These findings suggest that while academic research emphasizes expanding task capabilities, industry adoption is constrained by performance, reliability, and safety requirements, underscoring the need for closer alignment between research and practical application.
@@
  • We identified and summarized nine types of tasks undertaken by LLMs in UAV systems and compared their distribution in academic literature and industrial projects. The results reveal the current applications and differences of UAV-LLM integration in academia and industry, providing a reference for future application-oriented research.
  • Through a practitioner survey, we analyzed their views on UAV-LLM integration. The findings reveal the potential reasons for the gap between academic and industrial efforts and emphasize the need to narrow this gap.
@@
  • We summarized challenges for future development, provided recommendations for researchers and practitioners, and released our experimental data and results at https://github.com/UAV LLMs/OnlineSupplementMaterial to support future research.
[…]
2. Background
2.1. Unmanned Aerial Vehicles
@@
As an aerial mobile platform with perception, decision-making, and execution capabilities, UAVs have been widely used in mission scenarios such as agricultural inspection (Tokekar et al., 2016), disaster response (Erdemir et al., 2020), logistics distribution (Stolaroff et al., 2018), and urban monitoring (Goodchild and Toy, 2018). With the diversification of application tasks, the dynamic operation environment, and the complexity of mission requirements, UAV systems are facing a growing demand for intelligence, including more sophisticated control strategies, more efficient path planning, and more natural human-computer interaction mechanisms (Shakhatreh et al., 2019).
@@
To support this evolution, open-source autopilot platforms and communication frameworks have become an important foundation of UAV research and development. PX4 (Meier et al., 2015) provides a widely used flight control software stack that supports both research and industrial UAV applications. MAVLink (Koubaa et al., 2019) offers a lightweight communication protocol that allows reliable message exchange between UAVs and ground stations. The ArduPilot project (Baldi et al., 2022), with extensive documentation and community support through the ArduPilot Wiki, enables developers to build and customize UAV systems for diverse mission profiles. Traditional rule-based control methods have gradually exposed their limitations in dealing with these challenges, driving the need for UAV technology to evolve to a higher level of intelligence.
2.2. Large Language Models
@@
Large language models (LLMs) have demonstrated strong generalization and semantic understanding capabilities in natural language processing, code generation, automatic question answering, reasoning planning, and other tasks (Naveed et al., 2023). Current mainstream LLMs (such as the GPT series, Claude, etc.) are based on the Transformer architecture. They can extract deep semantics from natural language, perform logical reasoning, and model tasks. Furthermore, LLMs have gradually acquired multimodal input processing capabilities, showing broad prospects for integration with embedded systems in the fields of visual language understanding, robot control, and task planning.
2.3. Applications of LLMs in UAV Systems
@@
The integration of UAVs with LLMs enables flight platforms to directly understand human instructions and carry them out, lowering the barrier to use and enabling more flexible operation. Some recent studies have already explored this possibility. For example, UAV-GPT (Koubaa et al., 2025) shows how a natural command like “please patrol this area” can be translated into structured parameters that the UAV then executes. In another direction, Grounded-SAM combined with GPT-4V (Ren et al., 2024; Limberg et al., 2024) demonstrates how language models can work with visual input to label aerial images, recognize targets, and generate natural descriptions.
@@
2.4. Card Sorting Method
@@
In order to functionally classify and structure the literature on UAVs and large language models, this paper adopts the card sorting method, which is a cognitive organization method widely used in information architecture design and classification system construction  (Wood and Wood, 2008; Spencer, 2009). In the card sorting process, researchers first create a card for each analysis object (such as a document), which contains key information related to the classification task, such as functional description, model type, or task instance. Then, researchers classify the cards with semantic associations in a bottom-up manner, thereby gradually building a logically consistent classification system. According to the degree of presetting of categories, the card sorting method can be divided into three types: open card sorting, closed card sorting, and hybrid card sorting. In closed card sorting, all categories are preset before sorting, and users only need to classify items into fixed categories. In open card sorting, categories are not preset, but are dynamically generated by participants according to data characteristics during the sorting process. Hybrid card sorting combines the characteristics of both, allowing expansion based on existing categories to achieve both systematicity and flexibility.
[…]
@@
3. RQ1: What is the current research status and focus of large language models in the field of UAVs?
@@
LLMs have powerful instruction understanding, task reasoning, and interactive control capabilities. Introducing LLMs into UAV systems is expected to achieve stronger intelligent perception and task coordination capabilities. However, most existing studies focus on specific models or application cases, and there is still a lack of systematic classification of the UAV-LLM integration. This not only limits the generalization application of new technologies, but also hinders the exploration of cross-domain solutions to key problems (such as task planning, information interaction, etc.). Therefore, it is necessary to comprehensively identify and summarize the current application fields and task categories through a systematic literature review, and identify future research directions in the UAV-LLM integration.
3.1. Systematic Literature Review (SLR)
@@
This study followed the systematic literature review (SLR) method guidelines proposed by Kitchenham et al. (Kitchenham et al., 2009) and carried out the research work according to its process. The specific process is shown in Figure 1, which includes three stages: literature search, literature selection, and data analysis. First, a literature search was conducted in databases including Google Scholar, ACM, and IEEE using a keyword-based strategy, and a total of 997 initial documents were obtained. Secondly, based on five exclusion criteria and combined with manual abstract analysis and snowball strategy (Bodin et al., 2019), 74 documents that were relevant to the research question were screened out. Finally, the selected documents were systematically read, and through card classification and inductive analysis, the answer to research question RQ1 was finally formed.
Figure 1. Methodological process for answering RQ1 via systematic literature review.
3.1.1. Literature Search
@@
To collect high-quality research literature on the theme of UAV-LLM integration, we first conducted a preliminary search using the Google Scholar engine and obtained 997 candidate papers. In our data collection strategy, we required that the paper title contain at least one of the keywords “Unmanned Aerial Vehicle (UAV)” or “Large Language Model (LLM)” to ensure broad coverage of the field. However, many works primarily discuss other aspects of UAV systems, such as hardware platform design, flight control engineering, or photogrammetry applications (Gupta et al., 2015; Shakhatreh et al., 2019; Colomina and Molina, 2014), which fall outside the scope of our focus on UAV-LLM integration. Therefore, we further filtered the corpus by requiring that the abstracts explicitly mention both “UAV” and “LLM”, ensuring that the final dataset directly relates to the integration of these two technologies.
3.1.2. Literature Selection

### v2

# Title for citation

@@
Abstract
In recent years, unmanned aerial vehicles (UAVs) have become increasingly popular in our daily lives and have attracted significant research interest in software engineering. At the same time, large language models (LLMs) have made notable advancements in language understanding, reasoning, and generation, making LLM applications in UAVs a promising research direction. However, existing studies have largely remained in preliminary exploration with a limited understanding of real-world practice, which causes an academia-industry gap and hinders the application of LLMs in UAVs. To address this, we conducted the first empirical study to investigate how LLMs support UAVs. To characterize common tasks and application scenarios of real-world UAV-LLM practices, we conducted a large-scale empirical study involving 997 research papers and 1,509 GitHub projects. The results classified nine common tasks (e.g., Natural Language Command Parsing) in four UAV workflows (e.g., Information Input) undertaken by LLMs in real-world UAV projects and revealed a large difference in the task distribution of research efforts and industry practices. To gain deeper insight into these differences and understand developers’ perspectives on the application of LLMs in UAVs, we conducted a survey of practitioners, receiving 52 valid responses from 15 countries. The results revealed that while 40.4% of developers have attempted to apply LLMs to UAV tasks, 59.6% still face challenges integrating their UAV projects with advanced LLM capabilities. Their feedback attributes these challenges to five factors, including technological maturity, performance, safety, cost, and others, and provides practical implications for researchers and developers in conducting UAV-LLM practices.
keywords:
@@ Abstract
1 Introduction
In recent years, unmanned aerial vehicles (UAVs) have become more and more popular in our daily lives and have played an increasingly important role [shakhatreh2019unmanned, liu2021precision, colomina2014unmanned]. They have attracted substantial research interest in software engineering, aiming to efficiently and securely implement advanced capabilities in UAV projects [wang2025dpfuzzer, khatiri2024simulation, minn2024dronlomaly, wang2025routhsearch]. However, while UAV operations involve complex processes of environmental perception, flight control, and task planning, current UAV systems still primarily rely on manual operation or rule-based pipelines [gupta2016survey], which limits their flexibility and adaptability in dynamic environments and diverse tasks. Consequently, it is still difficult to achieve scalable high-autonomy operations and thereby constraining the broader deployment of UAVs in more complex scenarios [colomina2014unmanned]. As UAV applications expand, these limitations become more noticeable, especially in situations that require quick reactions, varied tasks, or unexpected adjustments.
Advances in large language models (LLMs) in language understanding, reasoning, and generation [chang2024survey, zhao2023llmsurvey] offer new possibilities to address the above limitations. By enabling UAVs to incorporate reasoning and interpretation in complex environments, LLMs reduce reliance on manual operation and rule-based pipelines, allowing tasks to be decomposed and decisions to be made with greater flexibility [tian2025uavs, wang2024llmrobotics, zeng2023llmrobotics, ahn2022saycan, brohan2023rt2]. This capability can significantly enhance the autonomy and adaptability of UAVs, making LLM applications in UAVs an increasingly prominent research direction.
Although existing studies [javaid2024large, tian2025uavs, tazir2023words] demonstrate the feasibility of this direction, most remain confined to simulation or controlled environments, with limited coverage of scenarios and task diversity. They often give insufficient consideration to critical factors in real-world applications, such as computational resources, latency, reliability, and compliance [survey2023uavedge, wu2024safety], and the defined task metrics do not readily correspond to the complexity of practical demands. These limitations risk creating a mismatch between academic research and actual needs, thereby reducing the practical value of research studies in translating them into techniques usable in UAV projects.
@@ 1 Introduction
RQ1. What are the current research focuses and application scenarios of large language models in UAV systems?
@@
Identifying the specific tasks that LLMs can perform in UAV systems is the first step toward understanding the advantages and limitations of their integration. To this end, we conducted a systematic literature review (SLR) [kitchenham2009systematic] and analyzed a total of 997 research papers related to LLMs and UAVs, applying a card sorting method [spencer2009cardsorting] to summarize and classify their contents. Based on the SLR process, we classify the support of LLMs for UAV workflows into four core stages in the UAV workflows—Information Input, Task Planning, Control Execution, and Result Feedback—each corresponding to a set of LLM-enabled tasks. Among these research directions, the top three most frequently studied directions are Single-UAV Task Reasoning and Planning, Control Logic and Command Generation, and Natural Language Command Parsing. Our characterization of UAV-LLMs research clarifies how LLMs are currently applied within UAV projects, which also enables us to compare researchers’ work with developers’ practices in subsequent analysis.
@@
RQ2. What are the differences between researchers’ and developers’ practices in integrating UAV projects with large language models?
Due to differences in objectives, resources, and evaluation standards, researchers and developers often emphasize different aspects when using new techniques. In this study, we aim to compare these differences and explore how the two sides can be better connected. We collected 1,509 UAV-LLM-related projects from GitHub and performed an empirical analysis of their characteristics and development practices.
@@
Our analysis revealed different efforts between researchers and developers in the UAV-LLM field. Researchers focus more on the design of complex system integration and theoretical mechanisms. It is committed to exploring how to use LLMs to improve system autonomy, flexibility, and multi-tasking capabilities. In contrast, developers emphasize system stability, low-cost implementation, and rapid deployment, and generally adopt efficient solutions. This, to some extent, limits the in-depth application of LLMs in scenarios requiring high timeliness or operating under resource constraints. By revealing the differences between researchers’ and developers’ practices, our results could provide practical implications for facilitating real-world UAV-LLM practices.
@@
RQ3. How do real-world developers perceive these differences, and what factors contribute to them?
Although our comparative analysis reveals the academia-industry gap for LLM applications in UAVs, the root causes and solutions of these differences remain unclear. To this end, we designed a survey for real-world developers, focusing on the key obstacles and design choices they face when using LLMs for UAVs.
@@
The analysis of 52 real-world responses from 15 countries shows that the application of LLMs in UAVs remains at an early stage. Only 19.05% of teams have deployed such projects, while 59.6% have not attempted application, citing reasons such as existing technology has met the project requirements (54.8%), insufficient performance of current LLMs (51.6%), and security risks of large language models (45.2%). The assessed technical maturity of all tasks is below three on a 5-point scale, with relatively higher scores for Information Summarization and Report Generation and LLM-based Chatbot, but lower scores for Multi-UAV Task Reasoning and Planning and Control Logic and Command Generation. Furthermore, 82.7% of developers prefer Onboard Control (Internal) Modes over direct control, reflecting caution toward applying LLMs in safety-critical operations. These findings suggest that while researchers emphasize expanding task capabilities, developers’ adoption is constrained by performance, reliability, and safety requirements, underscoring the need for closer alignment between research and practical application.
@@
  • We identified and summarized nine types of tasks undertaken by LLMs in UAV projects and revealed the differences of their distributions in academic literature and industrial projects. The results reveal current applications and differences in LLM applications in UAVs among researchers and developers, providing a reference for future application-oriented research.
  • Through a large-scale real-world empirical study, we analyzed developers’ perspectives on LLM applications in UAVs. The findings identify potential reasons for the gap between researchers’ and developers’ practices and emphasize the need to bridge it.
@@
  • Our empirical results demonstrated challenges for future development and provided practical implications for researchers and developers. To support future research, our experimental data and analysis results are available at https://github.com/UAV LLMs/OnlineSupplementMaterial.
2 Background
2.1 Unmanned Aerial Vehicles
@@
As an aerial mobile platform with perception, decision-making, and execution capabilities, UAVs have been widely used in our daily lives [tokekar2016sensor, ho2022open, erdemir2020autonomous, stolaroff2018energy, goodchild2018delivery]. With the diversification of application tasks, the dynamic operation environment, and the complexity of mission requirements, UAV systems are facing a growing demand for intelligence, including more sophisticated control strategies, more efficient path planning, and more natural human-computer interaction mechanisms [shakhatreh2019unmanned].
@@
To support this evolution, open-source autopilot platforms and communication frameworks have become an important foundation of UAV research and development. PX4 [meier2015px4] provides a widely used flight control software stack that supports both research and industrial UAV applications. MAVLink [koubaa2019mavlink] offers a lightweight communication protocol that allows reliable message exchange between UAVs and ground stations. The ArduPilot project [baldi2022ardupilot], with extensive documentation and community support through the ArduPilot Wiki, enables developers to build and customize UAV systems for diverse mission profiles. Traditional rule-based control methods have gradually exposed their limitations in dealing with these challenges, driving the need for UAV techniques to evolve to a higher level of intelligence.
2.2 Large Language Models
@@
Large language models (LLMs) have demonstrated strong generalization and semantic understanding capabilities in natural language processing, code generation, automatic question answering, reasoning planning, and other tasks [naveed2023comprehensive]. Current mainstream LLMs (such as the ChatGPT [schulman2022chatgpt], Claude [anthropic2024claude35]) are based on the Transformer architecture [abadi2015tensorflow]. They can extract deep semantics from natural language, perform logical reasoning, and model tasks. Furthermore, LLMs have gradually acquired multimodal input processing capabilities, showing broad prospects for integration with embedded systems in the fields of visual language understanding, robot control, and task planning.
2.3 Applications of LLMs in UAV Systems
@@
The integration of UAVs with LLMs enables flight platforms to directly understand human instructions and carry them out, lowering the barrier to use and enabling more flexible operation. Some recent studies have already explored this possibility. For example, UAV-GPT [koubaa2025next] shows how a natural command like “please patrol this area” can be translated into structured parameters that the UAV then executes. In another direction, Grounded-SAM combined with GPT-4V [ren2024grounded, limberg2024leveraging] demonstrates how language models can work with visual input to label aerial images, recognize targets, and generate natural descriptions.
@@
2.4 Card Sorting Method
@@
In order to functionally classify and structure the literature on UAVs and large language models, this paper adopts the card sorting method, which is a cognitive organization method widely used in information architecture design and classification system construction  [wood2008card, spencer2009cardsorting]. In the card sorting process, researchers first create a card for each analysis object (e.g., a document) that contains key information related to the classification task, such as a functional description, model type, or task instance. Then, researchers classify the cards with semantic associations in a bottom-up manner, thereby gradually building a logically consistent classification system. According to the degree of presetting of categories, the card sorting method can be divided into three types: open card sorting, closed card sorting, and hybrid card sorting. In closed card sorting, all categories are preset before sorting, and users only need to classify items into fixed categories. In open card sorting, categories are not preset, but are dynamically generated by participants according to data characteristics during the sorting process. Hybrid card sorting combines the characteristics of both, allowing expansion based on existing categories to achieve both systematicity and flexibility.
@@
3 RQ1: What are the current research focuses and application scenarios of large language models in UAV systems?
@@
LLMs have powerful instruction understanding, task reasoning, and interactive control capabilities. Introducing LLMs into UAV systems is expected to achieve stronger intelligent perception and task coordination capabilities. However, most existing studies focus on specific models or application cases, and there is still a lack of systematic classification of the UAV-LLM integration. This not only limits the generalization application of new techniques, but also hinders the exploration of cross-domain solutions to key problems (e.g., task planning, information interaction). Therefore, it is necessary to comprehensively identify and summarize current application fields and task categories through a systematic literature review, and to identify future research directions in UAV-LLM integration.
3.1 Systematic Literature Review (SLR)
@@
This study followed the systematic literature review (SLR) method guidelines proposed by Kitchenham et al. [kitchenham2009systematic] and carried out the research work according to its process. The specific process is shown in Figure 1, which includes three stages: literature search, literature selection, and data analysis. First, a literature search was conducted in databases including Google Scholar, ACM, and IEEE using a keyword-based strategy, and a total of 997 initial documents were obtained. Secondly, based on five exclusion criteria and combined with manual abstract analysis and snowball strategy [naderifar2017snowball], 74 documents that were relevant to the research question were screened out. Finally, the selected documents were systematically read, and through card classification and inductive analysis, the answer to research question RQ1 was finally formed.
Figure 1: Methodological process for answering RQ1 via systematic literature review.
3.1.1 Literature Search
@@
To collect high-quality research literature on the theme of UAV-LLM integration, we first conducted a preliminary search using the Google Scholar engine and obtained 997 candidate papers. In our data collection strategy, we required that the paper title contain at least one of the keywords “Unmanned Aerial Vehicle (UAV)” or “Large Language Model (LLM)” to ensure broad coverage of the field. However, many works primarily discuss other aspects of UAV systems, such as hardware platform design, flight control engineering, or photogrammetry applications [ahmed2022survey, shakhatreh2019unmanned, colomina2014unmanned], which fall outside the scope of our focus on UAV-LLM integration. Therefore, we further filtered the corpus by requiring that the abstracts explicitly mention both “UAV” and “LLM”, ensuring that the final dataset directly relates to the integration of these two.
3.1.2 Literature Selection